OpenStudy (anonymous):

If you save $4,500 at an interest rate of 13 percent per year, how much will you have at the end of seven years?

OpenStudy (kropot72):

Do you know whether the interest is compounded yearly or is it simple interest?

OpenStudy (abmon98):

Simple Interest Simple interest is charged only on the principal amount. The following formula can be used to calculate simple interest: Simple Interest (Is) = P × i × t Is=4500*13%*7